# Answer, Refuse, or Guess? Investigating Risk-Aware Decision Making in Language Models

### v1

@@ A.2 Prompt Variations for Main Experiments
In our main experiment results (see section 4.4), we report the averaged performance from three prompt variations to enhance the robustness of our findings (see section 4.2). Here we list the other two prompt variations in Figure 15, 16, 17, 18, 19, and 20 except the ones shown in the main text.
Appendix B Performance of Individual Prompts and Error Bars
@@
We show performance scores of three individual prompt variations used in the main experiments (section 4.4) in Table 5, 6, and 7. We also list the standard deviations of the three prompts in Table 8.
Appendix C Implementation Details
@@
For the six LMs used in Table 2, including gpt-4o-2024-08-06, gpt-4o-mini-2024-07-18, claude-3-5-sonnet-20241022, claude-3-5-haiku-20241022, gemini-1.5-pro-002, and gemini-1.5-flash-002, we use greedy decoding (temperature = 0) to make the output as deterministic as possible for better reproducibility. For the reasoning LMs used in Table 3, we use the default decoding temperatures suggested by the model developers. The temperatures are 1.0 for o3-mini-2025-01-31 and 0.6 for DeepSeek-R1. As for the maximum number of output tokens, we set to 4,096 tokens for non-reasoning LMs, 25,000 tokens for o3-mini-2025-01-31, and 32,768 tokens for DeepSeek-R1. We manually inspect LMs’ output and make sure that their output won’t be truncated under such maximum token constraints.
@@
Appendix D Why Stepwise Prompts Perform Worse than Prompt Chaining at High-Risk?
From the main results in Table 2, we find that stepwise prompts generally perform worse than prompt chaining in the high-risk settings. Since both methods (see Figure 6) use skill decomposition to solve the task, the final score $S$ (see section 2.1) depends on the performance of each individual skill. Therefore, we can inspect how well do LMs perform the individual skills by the following metrics:
@@ Appendix D Why Stepwise Prompts Perform Worse than Prompt Chaining at High-Risk?
    Expected-Value Reasoning: As shown by Figure 7 in section 5.1, LMs sometimes fail to explicitly calculate expected values. We thus measure this ability by the proportion of using expected-value reasoning (EVR).
We list metrics for these individual skills and the final performance $S$ in Table 9, which uses $(s_{\mathrm{cor}},s_{\mathrm{inc}})=(1,-8)$ as the representative risk level. From the results of “averaged across all models”, we can see that prompt chaining generally performs better at each individual skill, except for ECE in MedQA. Accordingly, it is probable that the superior performance of prompt chaining arises from the fact that it performs better at each skill.
Appendix E Better Confidence Estimation (Calibration) Improves Final Scores
As mentioned in the limitations (§Limitations), improving the individual skills has the potential to increase the overall performance of prompt chaining on risk-aware decision making. Here, we show evidence that improving confidence estimation (i.e., calibration) increases the final score in most cases.
@@ Appendix E Better Confidence Estimation (Calibration) Improves Final Scores
    Predicted answer = the answer letter only: We only include the answer letter by extracting the letter from “ANSWER: <letter>” in the LMs’ output from “Skill 1: Question Answering” (see the upper right prompt in Figure 6).
@@
The comparison between these two choices is illustrated in Figure 10. We show the calibration curves / metrics of the two choices in Figure 11, which reveals that including rationales in the predicted answer increases the expected calibration error (ECE). On the other hand, including only the answer letters leads to better calibration. We then compare the final scores of the two different choices of confidence estimation, and show the results in Table 4. Overall, the results indicate that the choice with better calibration performance indeed leads to better final performance in 10 out of 12 cases, so we use choice 1 in our main experiments (§4.4).
Appendix F License for Code and Datasets
@@
Table 8: Standard deviations of three prompt variations, of which the performance is listed in Table 5, 6, and 7.
@@
Metric ACC($\uparrow$) ECE($\downarrow$) EVR($\uparrow$) $S(\uparrow)$     ACC($\uparrow$) ECE($\downarrow$) EVR($\uparrow$) $S(\uparrow)$     ACC($\uparrow$) ECE($\downarrow$) EVR($\uparrow$) $S(\uparrow)$
@@
Figure 12: The first example of a paraphrased pure gambling prompt.
Figure 13: The second example of a paraphrased pure gambling prompt.
Figure 14: The third example of a paraphrased pure gambling prompt.
Figure 15: The second variation of the no risk informed prompt.
No risk informed prompt - variation 3 You are given a multiple-choice question. Please provide a detailed explanation of your reasoning before selecting the correct answer. Your final answer should be presented in the format: "ANSWER: <letter>", where <letter> corresponds to the chosen option. {question} {choices_text}
Figure 16: The third variation of the no risk informed prompt.
Risk-informing prompt - variation 2 Please respond to the question provided with the given options. Question: {question} Choices: {choices_text} The scoring criteria is as follows: {scoring_criteria} Consider the scoring criteria outlined below to achieve the highest possible score. You have the option to either select one of the provided choices by using the corresponding letter (A, B, C, or D) or decline to answer by choosing N. Begin with your reasoning, followed by your final choice in the format: ANSWER: <letter>
Figure 17: The second variation of the risk-informing prompt.
Risk-informing prompt - variation 3 Respond to the following question using the provided answer choices. Question: {question} Choices: {choices_text} The scoring criteria is as follows: {scoring_criteria} Keep in mind the scoring guidelines to maximize your score for this question. You may either provide your answer by selecting one of the letters A, B, C, or D, or choose not to answer by selecting N. Start with your reasoning, then conclude with your answer in the format: ANSWER: <letter>
Figure 18: The third variation of the risk-informing prompt.
Stepwise prompt - variation 2 Respond to the question provided with the given options. Question: {question} Choices: {choices_text} The scoring criteria is as follows: {scoring_criteria} Your objective is to achieve the highest possible score according to the above scoring criteria. To do this, please complete the following three tasks: 1. Provide your thought process and then give your final answer in the format: ANSWER: <letter> 2. Indicate your confidence in your predicted answer in the format: CONFIDENCE: <a score between 0% and 100%> 3. Decide whether to: a) Answer the question by providing your predicted answer with a single letter Y b) Decline to answer the question by outputting a single letter N Provide your reasoning first, followed by your final decision in the format: CHOICE: <letter>
Figure 19: The second variation of the stepwise prompt.
Stepwise prompt - variation 3 Please address the question below using the provided answer choices. Question: {question} Choices: {choices_text} The scoring guidelines are: {scoring_criteria} Your aim is to maximize your score for this question based on the scoring guidelines provided. To accomplish this, please adhere to the following three steps: 1. Offer your reasoning and then state your final answer in the format: ANSWER: <letter> 2. Specify your confidence level in your answer in the format: CONFIDENCE: <a score between 0% and 100%> 3. Choose to: a) Proceed with answering the question by outputting a single letter Y b) Opt out of answering the question by providing a single letter N Provide your reasoning first, followed by your final selection in the format: CHOICE: <letter>
Figure 20: The third variation of the stepwise prompt.

### v2

@@ B.3 Prompt Variations for Main Experiments
In our main experiment results (see section 4.2), we report the averaged performance from three prompt variations to enhance the robustness of our findings. Here we list the other two prompt variations in Figure 11, 12, 13, 14, 15, and 16 except the ones shown in the main text.
Figure 11: The second variation of the no risk informed prompt.
No risk informed prompt - variation 3 You are given a multiple-choice question. Please provide a detailed explanation of your reasoning before selecting the correct answer. Your final answer should be presented in the format: ”ANSWER: \$letter”, where \$letter corresponds to the chosen option. {question} {choices_text}
Figure 12: The third variation of the no risk informed prompt.
Risk-informing prompt - variation 2 Please respond to the question provided with the given options. Question: {question} Choices: {choices_text} The scoring criteria is as follows: {scoring_criteria} Consider the scoring criteria outlined below to achieve the highest possible score. You have the option to either select one of the provided choices by using the corresponding letter (A, B, C, or D) or decline to answer by choosing N. Begin with your reasoning, followed by your final choice in the format: ANSWER: ¡letter¿
Figure 13: The second variation of the risk-informing prompt.
Risk-informing prompt - variation 3 Respond to the following question using the provided answer choices. Question: {question} Choices: {choices_text} The scoring criteria is as follows: {scoring_criteria} Keep in mind the scoring guidelines to maximize your score for this question. You may either provide your answer by selecting one of the letters A, B, C, or D, or choose not to answer by selecting N. Start with your reasoning, then conclude with your answer in the format: ANSWER: ¡letter¿
Figure 14: The third variation of the risk-informing prompt.
Stepwise prompt - variation 2 Respond to the question provided with the given options. Question: {question} Choices: {choices_text} The scoring criteria is as follows: {scoring_criteria} Your objective is to achieve the highest possible score according to the above scoring criteria. To do this, please complete the following three tasks: 1. Provide your thought process and then give your final answer in the format: ANSWER: ¡letter¿ 2. Indicate your confidence in your predicted answer in the format: CONFIDENCE: ¡a score between 0% and 100%¿ 3. Decide whether to: a) Answer the question by providing your predicted answer with a single letter Y b) Decline to answer the question by outputting a single letter N Provide your reasoning first, followed by your final decision in the format: CHOICE: ¡letter¿
Figure 15: The second variation of the stepwise prompt.
Stepwise prompt - variation 3 Please address the question below using the provided answer choices. Question: {question} Choices: {choices_text} The scoring guidelines are: {scoring_criteria} Your aim is to maximize your score for this question based on the scoring guidelines provided. To accomplish this, please adhere to the following three steps: 1. Offer your reasoning and then state your final answer in the format: ANSWER: ¡letter¿ 2. Specify your confidence level in your answer in the format: CONFIDENCE: ¡a score between 0% and 100%¿ 3. Choose to: a) Proceed with answering the question by outputting a single letter Y b) Opt out of answering the question by providing a single letter N Provide your reasoning first, followed by your final selection in the format: CHOICE: ¡letter¿
Figure 16: The third variation of the stepwise prompt.
Appendix C Comparison between four methods in low-risk settings
Table 4 and Table 5 show empirical results of four methods in the low-risk settings. As expected, the no-risk prompt performs the best on average since $r_{\mathrm{guess}}>r_{\mathrm{ref}}=0$.
Table 4: Performance of four methods under $(r_{\mathrm{cor}},r_{\mathrm{inc}})=(4,-1)$
Method/Model C-Haiku GPT-mini Gem-Flash C-Sonnet GPT-4o Gem-Pro Average
Dataset MMLU
No risk informed 0.760 0.768 0.768 0.860 0.844 0.811 0.802
Risk-informing 0.754 0.769 0.767 0.852 0.836 0.814 0.799
Stepwise prompt 0.759 0.769 0.763 0.849 0.828 0.819 0.798
Prompt chaining 0.753 0.688 0.757 0.855 0.841 0.798 0.782
Dataset MedQA
No risk informed 0.714 0.747 0.667 0.868 0.882 0.768 0.774
Risk-informing 0.713 0.744 0.667 0.857 0.877 0.774 0.772
Stepwise prompt 0.717 0.743 0.668 0.858 0.877 0.778 0.773
Prompt chaining 0.712 0.659 0.611 0.863 0.880 0.723 0.741
Dataset GPQA
No risk informed 0.209 0.246 0.331 0.460 0.384 0.445 0.346
Risk-informing 0.217 0.254 0.329 0.434 0.362 0.452 0.341
Stepwise prompt 0.216 0.241 0.305 0.436 0.345 0.419 0.327
Prompt chaining 0.208 0.118 0.290 0.447 0.381 0.382 0.304
Table 5: Performance of four methods under $(r_{\mathrm{cor}},r_{\mathrm{inc}})=(8,-1)$
Method/Model C-Haiku GPT-mini Gem-Flash C-Sonnet GPT-4o Gem-Pro Average
Dataset MMLU
No risk informed 0.784 0.791 0.790 0.874 0.859 0.830 0.821
Risk-informing 0.777 0.787 0.792 0.866 0.850 0.834 0.818
Stepwise prompt 0.778 0.783 0.791 0.871 0.840 0.841 0.817
Prompt chaining 0.773 0.701 0.775 0.872 0.857 0.822 0.800
Dataset MedQA
No risk informed 0.742 0.772 0.700 0.881 0.893 0.791 0.797
Risk-informing 0.743 0.774 0.698 0.866 0.884 0.795 0.793
Stepwise prompt 0.738 0.765 0.703 0.874 0.891 0.801 0.795
Prompt chaining 0.735 0.677 0.643 0.878 0.893 0.745 0.762
Dataset GPQA
No risk informed 0.288 0.320 0.396 0.514 0.444 0.500 0.410
Risk-informing 0.298 0.323 0.369 0.504 0.424 0.484 0.400
Stepwise prompt 0.295 0.317 0.368 0.495 0.408 0.479 0.392
Prompt chaining 0.284 0.140 0.344 0.511 0.442 0.462 0.364
Appendix D Performance of Individual Prompts and Error Bars
@@
We show performance scores of three individual prompt variations used in the main experiments (section 4.2) in Table 7, 8, and 9. We also list the standard deviations of the three prompts in Table 10.
Appendix E Implementation Details
@@
For the six LMs used in Table 124, and 5, including gpt-4o-2024-08-06, gpt-4o-mini-2024-07-18, claude-3-5-sonnet-20241022, claude-3-5-haiku-20241022, gemini-1.5-pro-002, and gemini-1.5-flash-002, we use greedy decoding (temperature = 0) to make the output as deterministic as possible for better reproducibility. For the reasoning LMs used in Table 3, we use the default decoding temperatures suggested by the model developers. The temperatures are 1.0 for o3-mini-2025-01-31 and 0.6 for DeepSeek-R1. As for the maximum number of output tokens, we set to 4,096 tokens for non-reasoning LMs, 25,000 tokens for o3-mini-2025-01-31, and 32,768 tokens for DeepSeek-R1. We manually inspect LMs’ output and make sure that their output won’t be truncated under such maximum token constraints.
@@
Appendix F Why Stepwise Prompts Perform Worse than Prompt Chaining at High-Risk?
From the main results in Table 1 and 2, we find that stepwise prompts generally perform worse than prompt chaining in the high-risk settings. Since both methods (see Figure 6) use skill decomposition to solve the task, the final score $R$ depends on the performance of each individual skill. Therefore, we can inspect how well do LMs perform the individual skills by the following metrics:
@@ Appendix F Why Stepwise Prompts Perform Worse than Prompt Chaining at High-Risk?
    Expected-Value Reasoning: As shown by Figure 4, LMs sometimes fail to explicitly calculate expected values. We thus measure this ability by the proportion of using expected-value reasoning (EVR).
We list metrics for these individual skills and the final performance $R$ in Table 11, which uses $(r_{\mathrm{cor}},r_{\mathrm{inc}})=(1,-8)$ as the representative risk structure. From the results of “averaged across all models”, we can see that prompt chaining generally performs better at each individual skill, except for ECE in MedQA. Accordingly, it is probable that the superior performance of prompt chaining arises from the fact that it performs better at each skill.
Appendix G Better Confidence Estimation (Calibration) Improves Final Scores
As mentioned in the discussion (section 5), improving the individual skills has the potential to increase the overall performance of prompt chaining on risk-aware decision making. Here, we show evidence that improving confidence estimation (i.e., calibration) increases the final score in most cases.
@@ Appendix G Better Confidence Estimation (Calibration) Improves Final Scores
    Predicted answer = the answer letter only: We only include the answer letter by extracting the letter from “ANSWER: \$letter” in the LMs’ output from “Skill 1: Question Answering” (see the upper right prompt in Figure 6).
@@
The comparison between these two choices is illustrated in Figure 17. We show the calibration curves / metrics of the two choices in Figure 18, which reveals that including rationales in the predicted answer increases the expected calibration error (ECE). On the other hand, including only the answer letters leads to better calibration. We then compare the final scores of the two different choices of confidence estimation, and show the results in Table 6. Overall, the results indicate that the choice with better calibration performance indeed leads to better final performance in 10 out of 12 cases, so we use choice 1 in our main experiments (section 4.2).
Appendix H License for Code and Datasets
@@
Table 10: Standard deviations of three prompt variations, of which the performance is listed in Table 7, 8, and 9.
@@
Metric ACC($\uparrow$) ECE($\downarrow$) EVR($\uparrow$) $R(\uparrow)$     ACC($\uparrow$) ECE($\downarrow$) EVR($\uparrow$) $R(\uparrow)$     ACC($\uparrow$) ECE($\downarrow$) EVR($\uparrow$) $R(\uparrow)$